OpenStudy (anonymous):

What is the solution of the equation? Eliminate any extraneous solutions. Square root 3x+28-8

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt{3x+28}=8 }\) like this?

OpenStudy (anonymous):

Square root of 3x+28-8=x

OpenStudy (anonymous):

A. -9 and -4 B. 9 and -4 C. -9 D. -4

OpenStudy (sloppycanada):

First step - simplify. Assuming your solving for x.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt{3x+28}-8=x}\) right?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

ok, 1) add 8 to both sides 2) raise both sides to the second power.

OpenStudy (anonymous):

28 and 8 right?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sqrt{3x+28}-8\color{blue}{\bf +8}=x\color{blue}{\bf +8}}\)

OpenStudy (solomonzelman):

see what I did ?